ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60, 62 Taraskin, A. F., Samara Aircpace University, Russia On problems of statistics for SDE We shall study inference problems for stochastic processes which to serve as solutions of the different stochastic differential equations (SDE). Likelihood method are developed for inference about hypotheses that determine the “coefficients” of SDE. The data are supposed to be a continuously observed sample path. We are able to non-Gaussian limit distributions in asymptotic problems of the statistics examine. In particular, we present conditions which provide the consistence and also the convergence to non-Gaussian limit distribution of the maximum likelihood estimator.